# Fast and Accurate Decoder for the XZZX code Using Simulated Annealing

### v1

@@ IV.2.1 The MWPM decoding algorithm
We briefly review the algorithm of the MWPM decoder for the XZZX code [9].
The MWPM decoder is formulated as an algorithm on the decoding graph $(\mathcal{V}_{\bullet},\mathcal{E}_{\bullet})$, introduced in Sec. III. For $u,v\in S_{\bullet}$, we define the weighted distance $w(u,v)$ between the vertices $u=(u_{x},u_{y})$ and $v=(v_{x},v_{y})$ as the shorter of the following two path lengths:
  • The weighted Manhattan distance
    $\alpha_{z}\lvert u_{x}-v_{x}\rvert\,+\,\alpha_{x}\lvert u_{y}-v_{y}\rvert$, (23)
  • The length of ‘path via a virtual boundary vertex’ [35]. For $u,v\in S_{\mathrm{gray}}$, this path consists of only horizontal edges, which connect $u$ and the left or right boundary and connect $v$ and the left or right boundary. The length of the path equals to
[…]
    For $u,v\in S_{\mathrm{white}}$, this path consists of only vertical edges, which connect $u$ and the top or bottom boundary and connect $v$ and the top or bottom boundary. The length of the path equals to
[…]
We present the algorithm of the MWPM decoder in Algorithm 4. In Line 4, if the number of vertices $|S|$ is odd, to enable a perfect matching, the virtual boundary vertex is appended to the syndrome $S$ so that $|S|$ is an even number [35]. In Line 6, we create the list of all possible candidates for pairing
$K\coloneqq\{\text{unordered pair }(u,v)\,|\,u,v\in S_{\bullet},\,u\neq v\}.$ (26)
In Line 8, the MWPM_graph function finds a minimum-weight perfect matching (MWPM) [36], i.e., a perfect matching $L\subset K$ such that the total distance $\sum_{(u,v)\in L}w(u,v)$ is minimum. For our implementation, we employ the C++ library, Blossom V [37].
@@
1: $S_{\bullet}$ : Syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
2: Recovery chain $C_{\bullet}$ consistent with $S_{\bullet}$
@@
5:    Append ‘the virtual boundary vertex’ to $S_{\bullet}$
6:   $K\coloneqq\{\text{unordered pair }(u,v)\,|\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (26) 7:
@@
11:   for all $(u,v)\in L$ do
12:    Add edges which connect $u$ and $v$ to $C_{\bullet}$.
@@
1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Syndrome
2: Recovery configuration $C$ consistent with $S$
@@
In Line 12, for each matching pair $(u,v)\in L$, we connect the vertices $u$ and $v$ with a path which consists of edges, i.e., applying $Z$ ($X$) operators to the qubits located at the centers of horizontal (vertical) edges, respectively. The set of these edges forms a recovery chain $C_{\bullet}$.
The MWPM function invokes the MWPM_CHAIN function twice, i.e., for the syndromes $S_{\mathrm{white}}$ and $S_{\mathrm{gray}}$ on the white gray and faces, respectively. Note that, the syndromes on only faces of the same color may be connected.
The MWPM algorithm is too complicated to implement in a hardware system such as FPGA. Thus, we propose a simple variant algorithm in Sec. IV.3.
IV.2.2 Interpreting the MWPM decoder as an energy minimization problem
Let us interpret the MWPM decoder as energy minimization problem, then compare it with our SA decoder.
Since the perfect matching is equivalent to the syndrome consistency $\partial{C_{\bullet}}=S$, the total distance of such an error chain $C_{\bullet}$ is defined as
$w(C_{\bullet})\coloneqq\sum_{(u,v)\in L_{0}}w(u,v)$ (27)
where $L_{0}$ is a perfect matching. Then, the MWPM decoding algorithm is equivalent to find a recovery error chain $C_{\bullet}$:
$\arg\min_{C_{\bullet}}\left\{w(C_{\bullet})\,\middle|\,\partial{C_{\bullet}}=S\right\}.$ (28)
Since the weighted distance $w(u,v)$ is defined by Eqs. (23),(24) or (25), Eq. (27) is expressed as
$w(C_{\bullet})=\alpha_{z}\,n_{\mathrm{hor}}\,+\,\alpha_{x}\,n_{\mathrm{ver}}$ (29)
where $n_{\mathrm{hor}}$ ($n_{\mathrm{ver}}$) is the number of errors which reside on horizontal (vertical) edges in the error chain $C_{\bullet}$, respectively. In Eq. (29), the weight coefficients $\alpha_{x},\alpha_{z}$ can be calculated by Eq. (9) from the error probabilities $p_{x},p_{z}$ assuming we know.
Let us compare the total distance in Eq. (29) and the energy defined in Eq. (10). Since Eq. (29) does not contain the probability $p_{y}$, the MWPM decoder cannot properly consider $Y$ errors unless $p_{x}:p_{y}:p_{z}=1:2:1$, and therefore is not optimal. On the other hand, Eq. (10) contains the probability $p_{y}$, and hence our SA decoder is able to attain optimal accuracy. In Sec. V, we execute the MWPM decoder to evaluate the accuracy of our SA decoders for comparison.
@@ IV.3 The greedy matching decoder
As described in Sec. IV.2.1, the MWPM decoder utilizes the MWPM graph algorithm, i.e., the perfect matching giving the minimum of the total distance. Exchanging this ‘global’ minimality for a low computational cost and simple algorithm yields a greedy version of the perfect matching algorithm. In this subsection, we elaborate on our greedy matching algorithm. We employ this algorithm to construct a pure error $T(S)$, set as an initial configuration for our SA decoder in Algorithm 1.
We give our algorithm in Algorithm 5. In Line 10, the GREEDY_MATCHING_CHAIN function matches the closest pair $(u,v)\in K$ with respect to $w(u,v)$, and connects the vertices $u$ and $v$. In Line 12, it excludes all the pairs whose vertex is $u$ or $v$ from the candidate list $K$. This algorithm thus enables us to monotonically reduce the size of the list $K$.
We efficiently implement $K$ using priority queue [38, §6.5]. In the priority queue with $n$ elements (e.g., a binary heap), each of insertion and extract-min operations takes $\mathcal{O}(\log n)$; therefore, performing the operations for $n$ elements has time complexity $\mathcal{O}(n\log n)$. For our greedy matching decoder, the size of the priority queue is $n=\mathcal{O}(|S_{\bullet}|^{2})$, where $|S_{\bullet}|$ denotes the number of incorrect syndromes. Thus, the time complexity of our greedy matching decoder is $\mathcal{O}(|S_{\bullet}|^{2}\log{|S_{\bullet}|})$.
@@
1: $S_{\bullet}$ : Syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
2: Recovery chain $C_{\bullet}$ consistent with $S_{\bullet}$
@@
5:    Append ‘the virtual boundary vertex’ to $S_{\bullet}$
6:   $K\coloneqq\{\text{unordered pair }(u,v)\,|\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (26)
@@
10:    Pick a pair $(u,v)\in K$ such that $w(u,v)$ is minimum.
11:    Add edges which connect $u$ and $v$ to $C_{\bullet}$.
12:    Exclude all the pairs whose vertex is $u$ or $v$ from the list $K$.
@@
1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Syndrome
2: Recovery configuration $C$ consistent with $S$
@@
As well as the MWPM decoder, since our greedy matching decoder employs the weighted distance $w(u,v)$ which is calculated from the probabilities $p_{z}$ and $p_{x}$, it cannot take into account the probability $p_{y}$.
IV.4 Our SA decoder which starts with different pure errors
@@
To improve efficiency, rather than using an identical pure error repeatedly, it is preferable to use a different pure error for each SA run. To this end, we modify our greedy matching algorithm. Specifically, when multiple pairs of vertices have the same distance, their matching order is randomized using pseudo-random numbers. This modified algorithm is presented in Algorithm 6. In Algorithm 6, we create the candidate list $K_{w}$, defined as the set of pairs which give the distance $w$:
$K_{w}\coloneqq\{(u,v)\in K\,|\,w(u,v)=w\}.$ (30)
This newly introduced randomness may produce a different pure error for every invocation, while still remaining consistent with the given syndrome $S$. However, such pure errors differ not only in their stabilizer components but also in their logical components in the decomposition of Eq. (13). Therefore, we need to transform a logical operator to the one based on a ‘standard’ pure error, as follows.
We present SA decoder which utilizes Algorithm 6 in Algorithm 7. In Line 5, for the first SA, we construct a pure error $T_{1}(S)$, which is used as ‘standard’.
Let us denote by $T(S)$ a pure error constructed from the second SA onward. By applying Eq. (13), $T_{1}(S)$ can be decomposed as
$T_{1}(S)=T(S)\,L_{Q}\,G^{\prime}$ (31)
for $Q\in\{I,X,Z,Y\}$ and $G^{\prime}\in\mathcal{G}$. The logical operator $L_{Q}$ expresses ‘difference’ of equivalence classes of logical operators based on $T(S)$ and $T_{1}(S)$, and is obtained by $L_{Q}]=[T(S)\,T_{1}(S)]$.
From Eq. (31), it follows that
$\operatorname{Rep}_{T(S)}(L_{P})=\operatorname{Rep}_{T_{1}(S)}(L_{QP}),$
which implies that “the logical operator $L_{P}$ based on T(S)T(S)” equals to “the logical operator $L_{QP}$ based on $T_{1}(S)$”. Thus, for each $P\in\{I,X,Z,Y\}$, we store the calculated energy in $E^{(i)}_{L_{QP}}$ in Line 16.
It should be noted that, in the for loops in Lines 11 and 14 in Algorithm 7, each iteration is an independent task, and hence these tasks can obviously be parallelized. In addition, Line 13 can be parallelized with Lines 1516.
Algorithm 6 The greedy matching decoder which may produce a different pure error
1: $S_{\bullet}$ : Syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
2: Recovery chain $C_{\bullet}$ consistent with $S_{\bullet}$
@@
5:    Append ‘the virtual boundary vertex’ to $S_{\bullet}$
6:   $K\coloneqq\{\text{unordered pair }(u,v)\,|\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (26)
@@
10:    $w\coloneqq\min\{w(u,v)\,|\,(u,v)\in K\}$
11:    $K_{w}\coloneqq\{(u,v)\in K\,|\,w(u,v)=w\}$ $\triangleright$ Eq. (30)
12:    while $K_{w}$ is not empty do
13:      Pick a pair $(u,v)\in K_{w}$ randomly.
14:      Add edges which connect $u$ and $v$ to $C_{\bullet}$.
15:      Exclude all the pairs whose vertex is $u$ or $v$ from the lists $K_{w}$ and $K$.
@@
1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Syndrome
2: Recovery configuration $C$ consistent with $S$
@@
Algorithm 7 Decoder using simulated annealing which may start with different pure errors
1: $S$ : Syndrome for error configuration
2: Recovery configuration $C$
3: function SA_Decoder_Different_pure_error($S$)
4:   $\triangleright$  First SA   $\triangleleft$
5:   $T_{1}\leftarrow$ greedy_matching_different($S$)
6:   for all $P\in\{I,X,Z,Y\}$ do
7:    $C^{(1)}_{L_{P}}\coloneqq T_{1}\,L_{P}$ $\triangleright$ Initial configuration for SA
@@
9:
10:   $\triangleright$  Second and later SAs   $\triangleleft$
11:   for $i=2$ to $N_{\mathrm{SA}}$ do
12:    $T\leftarrow$ greedy_matching_different($S$)
13:    Determine $Q\in\{I,X,Z,Y\}$ such that $L_{Q}]=[T\,T_{1}]$.
14:    for all $P\in\{I,X,Z,Y\}$ do
15:      $C_{L_{P}}\coloneqq T\,L_{\mathrm{P}}$ $\triangleright$ Initial configuration for SA
16:      $E^{(i)}_{L_{QP}}\leftarrow$ simulated_annealing($C_{L_{P}}$)
17:
18:   for all $P\in\{I,X,Z,Y\}$ do
19:    $E_{L_{P}}\coloneqq\min\limits_{i=1,\dots,N_{\mathrm{SA}}}{E^{(i)}_{L_{P}}}$
20:
21:   $P^{*}\coloneqq\operatorname*{\arg\min}\limits_{P\in\{I,X,Y,Z\}}{E_{L_{P}}}$ $\triangleright$ Eq. (20) 22: 23:
24:   return $C^{(1)}_{L_{P^{*}}}$
Like the MWPM decoder, our greedy matching decoder is MLE-type; the MLD variant is obtained by disabling SA (i.e., by setting $N_{\beta}=0$) in Algorithm 7.
V Numerical result

### v2

@@ IV.2.1 The MWPM decoding algorithm
We briefly review the algorithm of the MWPM decoder for the XZZX code [xzzx_nature].
The MWPM decoder is formulated on the decoding graph $(\mathcal{V}_{\bullet},\mathcal{E}_{\bullet})$, introduced in Sec. III. We index the vertices by integer coordinates as
$\mathcal{V}_{\mathrm{white}}$ $=\{(i,j)\mid i\in\{1,\dots,d\},\;j\in\{1,\dots,d-1\}\},$
$\mathcal{V}_{\mathrm{gray}}$ $=\{(i,j)\mid i\in\{1,\dots,d-1\},\;j\in\{1,\dots,d\}\}$
where $i$ and $j$ denote the horizontal and vertical coordinates, respectively, and $(1,1)$ is the lower-left corner.
For vertices $u,v\in S_{\bullet}$, we define the weight $w(u,v)$ between $u=(u_{x},u_{y})$ and $v=(v_{x},v_{y})$
$w(u,v)\coloneqq\min\left\{w_{\mathrm{int}}(u,v),\,w_{\mathrm{bnd}}(u,v)\right\}.$ (25)
Here, $w_{\mathrm{int}}(u,v)$ is the weighted Manhattan distance [xzzx_nature]
$w_{\mathrm{int}}(u,v)\coloneqq\tilde{\alpha}_{z}\lvert u_{x}-v_{x}\rvert\,+\,\tilde{\alpha}_{x}\lvert u_{y}-v_{y}\rvert$, (26)
with
$\tilde{\alpha}_{\mu}\coloneqq-\log\!{\left(\frac{\tilde{p}_{\mu}}{1-p}\right)}\quad\text{for $\mu\in\{x,z\}$}$ (27)
where we defined the probabilities $\tilde{p}_{x}\coloneqq p_{x}+p_{y}$ and $\tilde{p}_{z}\coloneqq p_{z}+p_{y}$. Note that MWPM is invariant under multiplying all edge weights by a positive constant. Therefore, we may freely rescale the weights without changing the matching solution.
The alternative weight $w_{\mathrm{bnd}}(u,v)$ is the weight of connecting both vertices to the nearest boundary [MWPM_complexity]:
  • For $u,v\in S_{\mathrm{white}}$, we define
    $\begin{split}w_{\mathrm{bnd}}(u,v)\coloneqq\tilde{\alpha}_{x}(&\min\{\operatorname{dist}_{\mathrm{B}}(u),\operatorname{dist}_{\mathrm{T}}(u)\}\,+\\ &\min\{\operatorname{dist}_{\mathrm{B}}(v),\operatorname{dist}_{\mathrm{T}}(v)\}\,),\end{split}$ (28)
    where $\operatorname{dist}_{\mathrm{B}}(u)\coloneqq u_{y}$ and $\operatorname{dist}_{\mathrm{T}}(u)\coloneqq d-u_{y}$ are the vertical distances from the vertex $u$ to the bottom and top boundaries, respectively, and similarly for $v$.
  • For $u,v\in S_{\mathrm{gray}}$, we define
    $\begin{split}w_{\mathrm{bnd}}(u,v)\coloneqq\tilde{\alpha}_{z}(&\min\{\operatorname{dist}_{\mathrm{L}}(u),\operatorname{dist}_{\mathrm{R}}(u)\}\,+\\ &\min\{\operatorname{dist}_{\mathrm{L}}(v),\operatorname{dist}_{\mathrm{R}}(v)\}\,),\end{split}$ (29)
    where $\operatorname{dist}_{\mathrm{L}}(u)\coloneqq u_{x}$ and $\operatorname{dist}_{\mathrm{R}}(u)\coloneqq d-u_{x}$ are the horizontal distances from the vertex $u$ to the left and right boundaries, respectively, and similarly for $v$.
We present the algorithm of the MWPM decoder in Algorithm 4. In Line 4, if the number of vertices $|S_{\bullet}|$ is odd, a virtual boundary vertex $b$ is appended to the syndrome $S_{\bullet}$ so that a perfect matching exists (i.e., $|S_{\bullet}|$ becomes even) [MWPM_complexity]. In this case, we extend the weight function $w(\cdot,\cdot)$ by defining, for $u\in S_{\bullet}$,
$w(u,b)\coloneqq\begin{cases}\tilde{\alpha}_{x}\min\{\operatorname{dist}_{\mathrm{B}}(u),\operatorname{dist}_{\mathrm{T}}(u)\},&\text{for }\bullet=\mathrm{white},\\ \tilde{\alpha}_{z}\min\{\operatorname{dist}_{\mathrm{L}}(u),\operatorname{dist}_{\mathrm{R}}(u)\},&\text{for }\bullet=\mathrm{gray}.\end{cases}$ (30)
Note that $w_{\mathrm{bnd}}(u,v)=w(u,b)+w(v,b)$, corresponding to pairing both the defects $u,v$ with the boundary $b$.
In Line 6, we create the set of (unordered) pairs
$K\coloneqq\{\{u,v\}\,\mid\,u,v\in S_{\bullet},\,u\neq v\},$ (31)
which are all possible candidates for pairing. In Line 8, the MWPM_graph function finds a minimum-weight perfect matching (MWPM) [Edmonds], i.e., a perfect matching $L\subset K$ such that the total distance $\sum_{\{u,v\}\in L}w(u,v)$ is minimum. For our implementation, we employ the C++ library, Blossom V [BlossomV].
@@
1: $S_{\bullet}$ : Measured syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
2: $C_{\bullet}$ : Recovery chain, i.e., $\partial{C_{\bullet}}=S_{\bullet}$
@@
5:    Append the virtual boundary vertex $b$ to $S_{\bullet}$
6:   $K\coloneqq\{\{u,v\}\,\mid\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (31) 7:
@@
11:   for all $\{u,v\}\in L$ do
12:    Add to $C_{\bullet}$ the edges along a path connecting $u$ and $v$ with the weight $w(u,v)$.
@@
1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Measured syndrome
2: $C$ : Recovery configuration, i.e., $\partial{C}=S$
@@
In Line 12, for each matched pair $\{u,v\}\in L$, we connect the vertices $u$ and $v$ with a path which consists of edges, i.e., applying $Z$ and $X$ operators to the qubits located at the centers of horizontal and vertical edges, respectively. This path has weight $w(u,v)$, i.e., it is chosen to attain the distance defined by Eqs. (26), (28), (29), or (30). The set of these edges forms a recovery chain $C_{\bullet}$.
The MWPM function invokes the MWPM_chain function twice, i.e., for the syndromes $S_{\mathrm{white}}$ and $S_{\mathrm{gray}}$ on the white and gray faces, respectively. Note that only syndromes on faces of the same color may be connected.
Implementing the MWPM decoder efficiently on hardware such as FPGAs is challenging due to its algorithmic complexity and memory/communication requirements. Thus, we propose a simple variant algorithm in Sec. IV.3.
IV.2.2 The MWPM decoder vs. our SA decoder as a minimization problem
Given a perfect matching $L\subset K$, we define its total weight as
$w(L)\coloneqq\sum_{\{u,v\}\in L}w(u,v).$ (32)
The MWPM decoder finds a minimum-weight perfect matching $L^{\ast}$:
$L^{\ast}\in\arg\min_{L}\left\{w(L)\,\middle|\,L\text{ is a perfect matching on }S_{\bullet}\right\}.$
From a perfect matching $L$, we can construct a recovery chain $C_{\bullet}(L)$, i.e., $\partial{C_{\bullet}(L)}=S_{\bullet}$, by connecting each matched pair $\{u,v\}\in L$ with a path of weight w(u,v)w(u,v) [^5] Note that $C_{\bullet}(L)$ is not unique in general, due to degeneracies in the minimum-weight paths.. Since the weighted distance $w(u,v)$ is defined by Eqs. (26),(28), (29), or (30), Eq. (32) is expressed as
$w(L)=\tilde{\alpha}_{z}\,n_{\mathrm{hor}}\,+\,\tilde{\alpha}_{x}\,n_{\mathrm{ver}}$ (33)
where $n_{\mathrm{hor}}$ and $n_{\mathrm{ver}}$ are the numbers of occupied horizontal and vertical edges in the error chain $C_{\bullet}(L)$, i.e., the numbers of $Z$- and $X$-type components, respectively. A $Y$ error contributes to both $n_{\mathrm{hor}}$ and $n_{\mathrm{ver}}$.
We now compare the total distance in Eq. (33) and the energy defined in Eq. (10). In the MWPM approach for the XZZX code, one solves two separate matching problems on the white and gray decoding graphs. This effectively decouples the inference of the $X$- and $Z$-type components and does not model the fact that a $Y$ error simultaneously contributes to both components on the same data qubit; hence it cannot capture $Y$-induced correlations. Consequently, when $p_{y}$ is non-negligible, this MWPM approach is expected to be suboptimal compared to our SA decoder, which explicitly incorporates $p_{y}$ via the $Y$-error weight (through $\alpha_{y}$) in the energy $\mathcal{H}(C)$ in Eq. (10).
@@ IV.3 The greedy matching decoder
As described in Sec. IV.2.1, the MWPM decoder utilizes the MWPM graph algorithm, i.e., the perfect matching giving the minimum of the total distance. By replacing the global minimization of MWPM with a local greedy matching rule, we obtain a simple and low computational cost decoder. We employ the greedy matching decoding algorithm to construct a reference configuration $R(S)$, set as an initial configuration for our SA decoder in Algorithm 1.
We give our greedy matching decoding algorithm in Algorithm 5. In Line 10, the greedy_matching_chain function matches the closest pair $\{u,v\}\in K$ with respect to $w(u,v)$, and connects the vertices $u$ and $v$. Line 12 removes all pairs that contain $u$ or $v$ from the candidate list $K$ of vertex pairs. This algorithm thus enables us to monotonically reduce the size of the list $K$.
We efficiently implement $K$ using a priority queue [algorithm_Cormen, §6.5]. In the priority queue with $m$ elements (e.g., a binary heap), each of insertion and extract-min operations takes $\mathcal{O}(\log m)$; therefore, performing the operations for $m$ elements has time complexity $\mathcal{O}(m\log m)$. For our greedy matching decoder, the size of the priority queue is $m=\mathcal{O}(|S_{\bullet}|^{2})$, where $|S_{\bullet}|$ denotes the number of syndrome defects. Thus, the worst-case time complexity is $\mathcal{O}(|S_{\bullet}|^{2}\log|S_{\bullet}|)$ for each color. Since $|S_{\bullet}|\leq|\mathcal{V}_{\bullet}|=\mathcal{O}(d^{2})$, this yields the worst-case bound $\mathcal{O}(d^{4}\log d)$.
@@
1: $S_{\bullet}$ : Measured syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
2: $C_{\bullet}$ : Recovery chain, i.e., $\partial{C_{\bullet}}=S_{\bullet}$
@@
5:    Append the virtual boundary vertex $b$ to $S_{\bullet}$
6:   $K\coloneqq\{\{u,v\}\,|\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (31)
@@
10:    Pick a pair $\{u,v\}\in K$ such that $w(u,v)$ is minimum.
11:    Add to $C_{\bullet}$ the edges along a path connecting $u$ and $v$ with the weight $w(u,v)$.
12:    Remove from $K$ all pairs $\{x,y\}$ such that $x\in\{u,v\}$ or $y\in\{u,v\}$.
@@
1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Measured syndrome
2: $C$ : Recovery configuration, i.e., $\partial{C}=S$
@@
Like the MWPM decoder, our greedy matching decoder employs the weighted distance $w(u,v)$, which is calculated from $p_{z}$ and $p_{x}$, and therefore cannot take $p_{y}$ into account.
IV.4 Our SA decoder which starts with different reference configurations
@@
To improve efficiency, rather than using an identical reference configuration repeatedly, it is preferable to use a different reference configuration for each SA run. To this end, we modify our greedy matching algorithm by randomizing the tie-breaking among equal-weight pairs. Specifically, we randomize tie-breaking among equal-weight pairs using pseudo-random numbers. This modified algorithm is presented in Algorithm 6. In Algorithm 6, we form the subset
$K_{\min}\coloneqq\{\{u,v\}\in K\,\mid\,w(u,v)=w_{\min}\},$ (34)
i.e., the set of candidate pairs that attain the current minimum weight $w_{\min}$.
This newly introduced randomness may produce a different reference configuration for every invocation, while still satisfying $\partial{C}=S$. However, two such reference configurations can differ by both a stabilizer component $G$ and a logical operator component $L_{P}$ in the decomposition of Eq. (14). Therefore, to compare energies across runs consistently, we relabel the logical operator cosets obtained in each run as those defined with respect to a fixed “standard” reference configuration.
We present a second SA decoder that utilizes Algorithm 6 in Algorithm 7. In Line 4, we construct a reference configuration $R_{1}(S)\in\Omega_{S}$, which we use as the fixed standard reference.
Let $R_{i}(S)\in\Omega_{S}$ ($i\geq 2$) denote the reference configuration produced in the $i$th run. Using Eq. (14), $R_{i}(S)$ can be decomposed as
$R_{i}(S)=R_{1}(S)\,L_{Q}\,G^{\prime}$ (35)
for some $Q\in\{I,X,Y,Z\}$ and some $G^{\prime}\in\mathcal{G}$. Since $\partial(R_{i}(S)\,R_{1}(S))=\emptyset$, the equivalence class $R_{i}(S)\,R_{1}(S)]\in\overline{\mathcal{P}}/\mathcal{G}$ is well-defined and we obtain $R_{i}(S)\,R_{1}(S)]=[L_{Q}]$. The Pauli operator $Q$ can be determined by checking its commutation relations with the logical operators $L_{X}$ and $L_{Z}$.
Since stabilizers commute with logical operators, Eq. (35) implies the correspondence
$\operatorname{\mathcal{C}}_{R_{i}(S)}(L_{P})=\operatorname{\mathcal{C}}_{R_{1}(S)}(L_{QP}).$ (36)
where $QP$ denotes the Pauli multiplication in $\{I,X,Y,Z\}$ (up to global phase), consistent with our convention of ignoring global phases. Eq. (36) means that “the coset of $L_{P}$ with respect to $R_{i}(S)$” equals “the coset of $L_{QP}$ with respect to $R_{1}(S)$”. Thus, for each $P\in\{I,X,Y,Z\}$, we store the calculated energy in $E^{(i)}_{L_{QP}}$ in Line 14.
Note that, in the for loops in Lines 9 and 12 in Algorithm 7, each iteration is an independent task, and hence these tasks can trivially be parallelized. In addition, Line 11 can be parallelized with Lines 1314.
Algorithm 6 The greedy matching decoder which may produce a different reference configuration
1: $S_{\bullet}$ : Measured syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
2: $C_{\bullet}$ : Recovery chain, i.e., $\partial{C_{\bullet}}=S_{\bullet}$
@@
5:    Append the virtual boundary vertex $b$ to $S_{\bullet}$
6:   $K\coloneqq\{\{u,v\}\,|\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (31)
@@
10:    $w_{\min}\coloneqq\min\{w(u,v)\,|\,\{u,v\}\in K\}$
11:    $K_{\min}\coloneqq\{\{u,v\}\in K\,|\,w(u,v)=w_{\min}\}$ $\triangleright$ Eq. (34)
12:    while $K_{\min}$ is not empty do
13:      Pick a pair $\{u,v\}\in K_{\min}$ uniformly at random.
14:      Add to $C_{\bullet}$ the edges along a path connecting $u$ and $v$ with the weight $w(u,v)$.
15:      Remove from $K$ and $K_{\min}$ all pairs $\{x,y\}$ such that $x\in\{u,v\}$ or $y\in\{u,v\}$.
@@
1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Measured syndrome
2: $C$ : Recovery configuration, i.e., $\partial{C}=S$
@@
Algorithm 7 Decoder using simulated annealing which may start with different reference configurations
1: $S$ : Measured syndrome
2: $C$ : Recovery configuration, i.e., $\partial{C}=S$
3: function SA_decoder_different_reference($S$)
4:   $R_{1}\leftarrow$ greedy_matching_different($S$) $\triangleright$ standard reference configuration
5:   for all $P\in\{I,X,Y,Z\}$ do
6:    $C^{(1)}_{L_{P}}\coloneqq R_{1}\,L_{P}$ $\triangleright$ standard reference
@@
8:
9:   for $i=2$ to $N_{\mathrm{SA}}$ do
10:    $R_{i}\leftarrow$ greedy_matching_different($S$)
11:    Determine $Q\in\{I,X,Y,Z\}$ such that $L_{Q}]=[R_{i}\,R_{1}]$. $\triangleright$ e.g., via commutation with $L_{X}$ and $L_{Z}$
12:    for all $P\in\{I,X,Y,Z\}$ do
13:      $C^{(i)}_{L_{P}}\coloneqq R_{i}\,L_{P}$ $\triangleright$ Initial configuration for SA
[…]
15:
16:   for all $P\in\{I,X,Y,Z\}$ do
17:    $E_{L_{P}}\coloneqq\min\limits_{i=1,\dots,N_{\mathrm{SA}}}{E^{(i)}_{L_{P}}} \triangleright$ estimate of $E^{\min}_{L_{P}}$
18:
19:   $P^{\ast}\coloneqq\operatorname*{\arg\min}\limits_{P\in\{I,X,Y,Z\}}{E_{L_{P}}}$ $\triangleright$ Eq. (21) 20: 21:
22:   return $C^{(1)}_{L_{P^{\ast}}} \triangleright$ defined in Line 6
V Numerical results